OpenStudy (anonymous):

John is 45 years old and wants to retire at 65. He wishes to make monthly deposits in an account paying 12% compounded monthly so when he retires he can withdraw 280 dollars a month for 30 years. How much should John deposit each month?

hero (hero):

Working on it

OpenStudy (anonymous):

thank you!!!!!

hero (hero):

:)

hero (hero):

Still working on it...

hero (hero):

Not an easy one apparently

OpenStudy (anonymous):

no problem!!!

OpenStudy (anonymous):

hahah oh i know!

jimthompson5910 (jim_thompson5910):

This is a future value problem, but instead of finding the future value FV, they want the payment That formula is \[P = \frac{iFV}{(1+i)^n-1}\] So P=((0.12/12)*280*12*30)/((1+0.12/12)^(20*12)-1)=101.89 So he should deposit 101.89 a month (for 20 years)

OpenStudy (anonymous):

hmmmm that one didn't work

hero (hero):

How many chances do you get?

OpenStudy (anonymous):

three

hero (hero):

Okay

OpenStudy (anonymous):

got it!

jimthompson5910 (jim_thompson5910):

just curious, what did you get?

OpenStudy (anonymous):

well the numbers changed so it changed to John is 45 years old and wants to retire at 65. He wishes to make monthly deposits in an account paying 6% compounded monthly so when he retires he can withdraw 350 dollars a month for 30 years. How much should John deposit each month?

OpenStudy (anonymous):

i got 126.35

hero (hero):

What were your calculations?

jimthompson5910 (jim_thompson5910):

hmm must have used the wrong formula...

OpenStudy (anonymous):

v=350 {1-(1+.06/12)^-360/(.06/12)} v=58377.07 58377.07=P(1+.005)^240-1/(.005) P=126.35